The length of a rectangle is 9 cm more than its width. The area of the rectangle is 136 cm^2. What are its dimensions?
Answer by Cromlix(4381)

Width = x
Length = 9 + x
x(9 + x) = 136
9x + x^2 = 136
x^2 + 9x - 136 = 0
(x + 17)(x - 8) = 0
x + 17 = 0
x = -17 (no number x <0)
x - 8 = 0
x = 8
Width = 8cm
Length = 17cm
Hope this helps.
:-)